# Data Requirement Goal Modeling for Machine Learning Systems

### v1

@@ 5. Conclusion and Future work
We proposed the use of goal modeling for this task, as it provides an intuitive way to communicate with experts from different domains and analyze trade-offs. To construct the Data Requirements Goal Model (DRGM), we surveyed the literature to identify the challenges associated with developing data requirements for ML systems. These challenges were categorized under four themes and presented as soft goals. We then provided a customization mechanism to adapt the model to different contexts and ML problem types, based on insights from the surveyed gray and white literature.
The goal model and its customization mechanism were evaluated using two illustrative examples based on real-world problems. The first problem was a classification problem from the medical domain, and the second was a regression problem from the energy domain. After customizing the DRGM for each problem, the customized DRGM was evaluated using two dataset examples for each problem. The resulting evaluations aligned with the outcomes reported in the respective studies.
@@
Regarding limitations, the customization mechanism was designed specifically for the most common types of supervised ML problems, namely regression and classification. Additionally, some of the KPI values are derived from practitioners’ recommendations and gray literature. Users are also required to manually modify the DRGM using the provided flowcharts, which could lead to errors due to misunderstandings, especially among users unfamiliar with GRL. While we attempted to cover as many cases as possible, there remain domain-specific requirements, particularly concerning data legality and data representativeness, that may not be fully addressed. Thus, we do not claim that this model eliminates the need to consult legal or domain experts. However, the model and customization mechanism significantly minimize this need by streamlining the selection process and quickly filtering out inapplicable datasets. Moreover, this work is limited to traditional ML models and different KPIs, especially for data quantity, has to be addressed for deep learning models and LLM supported models.
@@
For future work, further evaluation of the model on a wider range of ML problems and contexts is recommended. Incorporating evaluations for pre-trained models is also necessary. Additionally, we plan to develop chatbot tool support to automate the customization mechanism, replacing the current manual process. Also, as this work is limited to traditional ML models, extra exploration is needed for deep learning models and LLM supported models.

### v2

@@ V Conclusion and Future work
The breakthroughs of ML in real-life applications highlighted the increasing need to address data requirements, as they are an essential building block of data-driven systems. This work aimed to guide non-experts in ML in identifying the essential data requirements based on the ML problem they are attempting to solve. We proposed the use of goal modeling for this task, as it provides an intuitive way to communicate with domain experts and to analyze trade-offs. To construct the Data Requirements Goal Model (DRGM), we surveyed the literature to identify the challenges associated with developing data requirements for ML systems. These challenges were categorized under four themes and presented as soft goals. We then provided a customization mechanism to adapt the model to different contexts and ML problem types, based on insights from the surveyed gray literature and scientific literature. The goal model and its customization mechanism were evaluated using two illustrative examples from two different contexts and ML problem types. The resulting evaluations aligned with the outcomes reported in the respective studies.
@@
Regarding limitations, the customization mechanism was designed for the most common types of supervised ML problems, namely regression and classification. In addition to the scientifically scholar publications, some of the KPI values are derived from practitioners’ recommendations and gray literature which may introduce subjectivity or lack the rigorous validation typically found in peer-reviewed sources. Users are also required to manually modify the DRGM using the provided flowcharts, which could lead to errors due to misunderstandings, especially among users unfamiliar with GRL. While we attempted to cover as many cases as possible, there remain domain-specific requirements, particularly concerning data legality and data representativeness, that may not be fully addressed. Thus, we do not claim that this model eliminates the need to consult legal or domain experts. However, the model and customization mechanism significantly minimize this need by streamlining the selection process and quickly filtering out inapplicable datasets. Moreover, this work is limited to traditional ML models ; different KPIs, especially related to data quantity, still need to be addressed for pre-trained deep learning models and LLM based applications.
@@
For future work, further evaluation of the model on a wider range of ML problems and contexts is recommended. Incorporating evaluations for pre-trained models is also necessary. Additionally, we plan to develop chatbot tool support to automate the customization mechanism, replacing the current manual process. Also, as this work is limited to traditional ML and deep learning models, extra exploration is needed for pre-trained deep learning models and LLM supported models.
ACKNOWLEDGMENTS
The authors acknowledge the support of King Fahd University of Petroleum and Minerals in the development of this work and acknowledge the support provided by the Deanship of Research at King Fahd University of Petroleum & Minerals (KFUPM).